OpenStudy (b2st):

2 = (√3b-2) - (√10-b) the answer is 6... HOW?! step by step please thanks!

OpenStudy (b2st):

\[2=\sqrt{3b-2} - \sqrt{10-b}\]

OpenStudy (b2st):

so this what i came up with? I'm not sure if I did it correctly (2)^2 = (√3b-2)^2 - √10-b (4)= 3b-2 (-√10-b) (4)^2= 3b-2 (-√10-b)^2 16=3b-2 +10-b 16 = 2b+8 12= 2b b= 6 ???????